40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Jan 30 19:26:03 2011, Koji, Summary, Green Locking, Prototype freq divider freq_divider.pdfIMG_3813.jpgIMG_3814.jpg
    Reply  Wed Feb 2 03:27:20 2011, Koji, Summary, Green Locking, 85MHz Freq divider freq_divider.pdfIMG_3816.JPGIMG_3818.JPG
       Reply  Wed Feb 2 09:56:55 2011, Koji, Summary, Green Locking, Installed the freq divider and Rana's PFD freq_divider_installation.pdf110201_freq_divider_output.pdf
          Reply  Wed Feb 2 10:44:26 2011, Koji, Summary, Green Locking, Freq fluctuation measured by the freq divider and Rana's analog PFD 110201_green_freq_fluctuation.pdf
             Reply  Wed Feb 2 12:55:34 2011, Koji, Summary, Green Locking, Freq fluctuation measured by the freq divider and Rana's analog PFD 
             Reply  Sat Feb 5 23:03:04 2011, rana, koji, Summary, Electronics, Analog Frequency Discriminator: splitter + mixer + long cable mixer.pdfUntitled.pngfsm.png
                Reply  Sun Feb 6 02:29:28 2011, rana, Update, Electronics, Analog MFD: longer cable mfd18.png
Message ID: 4239     Entry time: Wed Feb 2 10:44:26 2011     In reply to: 4238     Reply to this: 4240   4254
Author: Koji 
Type: Summary 
Category: Green Locking 
Subject: Freq fluctuation measured by the freq divider and Rana's analog PFD 

The freq fluctuation of the beat note has been measured with the following condition

  • The IR beam only locked to the MC. The green beam locked to the arm
  • Both of the IR and green locked to the x-arm

Calibration
- The output of the freq divider is already calibrated to have the unit of MHz.

- The transfer function between the analog PFD channel and the digital PFD output was measured to be -23dB = 0.7.
  The gain of the XARM-FINE channel was changed to 0.7 such that the output is calibrated in MHz.

Results

- I have not checked the analog noise level of the analog PFD path. We may need more whitening gain (by icreasing the gain of SR560).

- The analog PFD is always better than the digital PFD above 20Hz.

- Both the digital and analog PFD showed good agreement below 20Hz.
  Note the measurement was not simultaneous.

- When the arm is locked with the ETMX being actuated , the fluctuation of the arm length must be stabilized by a huge factor
(~10^5 according to Kiwamu's entry) However, we only could see the stabilization factor of 30.

As this residual is the difference of the freq noise felt by the IR and the green,
this is a real issue to be tackled.

- The RMS fluctuations of the arm with and without the IR beam being locked are 2MHz and 0.1MHz,
which correcponds to the arm length motion of 250nm and 13nm, respectively.
Ed: I had to use 532nm in stead of 1064nm. The correct numbers are 130nm and 7nm.

- Without the IR locked, The typical peak-to-peak fluctuation of the beat freq was 10MHz.

Attachment 1: 110201_green_freq_fluctuation.pdf  122 kB  | Hide | Hide all
110201_green_freq_fluctuation.pdf
ELOG V3.1.3-